# Multi-Level Embedding and Alignment Network with Consistency and Invariance Learning for Cross-View Geo-Localization

### v2

@@ III-D Cross-Domain Enhanced Alignment branch
In cross-view geo-localization tasks, images taken from different platforms often exhibit significant geometric differences, resulting in substantial variations in feature distributions and posing challenges for cross-domain matching. Feature alignment has been proven to be effective in addressing this issue[31, 32, 33]. Inspired by [19], we designed a cross-domain enhanced alignment (CEA) branch that uses multi-level fusion and adaptive calibration strategies to dynamically adjust feature consistency within a shared latent space in high-dimensional embeddings. In addition, invariance constraints are incorporated to learn the inherent structural invariances of these features. As shown in Fig.2 after the initial feature extraction by the backbone, we obtain high-dimensional features from different perspectives (drone view and satellite view ). To efficiently utilize global information and facilitate cross-view feature alignment and matching, we apply a transformation to flatten the spatial dimensions into a single dimension , obtaining reshaped features , which serve as the input features for the CEA branch. Here, , . Break down the subsequent procedures into three modules, including domain enhancement (DE) module, feature adaptive temperature (FAT) module, and multi-scale feature fusion (MSF) module, which are described as follows.
@@
DE module: We apply two convolutional kernels and , to project the two-dimensional input features into a higher-dimensional space. Subsequently, the batch normalization operation and the ReLU activation function are used to adjust the feature distribution, ensuring stability during the training phase. This procedure obtains a high-dimensional feature representation . Then, another convolutional kernel is used to map the high-dimensional features back to the low-dimensional space, followed by the dropout operation and the normalization operation to improve the generalization capability. This produces a more compact feature representation , which improves computational efficiency and can be formally expressed as follows,
@@
FAT module: After obtaining compact low-dimensional features , a 1×1 convolutional kernel is applied to further adjust the feature representation. We apply T-Softmax to achieve the temperature-scaled local features with temperature factor , highlighting the crucial distinctions of the features . This is followed by S-Softmax for global rebalancing to ensure that the features can maintain consistency and harmony across different levels or scales. We have
where the dual Softmax-processed feature . Both the input feature and the dual Softmax-processed features are transformed into a shared dimension of by the convolutional kernels and , respectively.
@@
MSF module: Once the features have been transformed into the same dimensional space, they are concatenated along the channel dimension, namely the concatenated features represented as follows,
@@
where . Then a convolutional kernel is applied to the concatenated features to restore the channel size back to , achieving a more compact representation. Finally, the batch normalization and the ReLU activation function are employed to normalize and activate the feature distributions. The output of the CEA branch is computed as follows,
@@ III-E Multi-Loss Optimization
For the feature embeddings and generated by the CEA branch described in Eq.(10), cosine similarity is used to estimate spatial directional consistency, while mean square errors measure absolute differences in the component components of the feature. Using cosine similarity and mean square errors, CDA loss effectively realizes the semantic consistency of aligned features at a global scale and maintains consistency in local details. The loss function can be formulated as follows,
where and are weighting coefficients to balance the contributions of cosine similarity and mean square errors . is designed to promote global semantic consistency and invariance across cross-view features and represented as follows,
where and , are denoted index of the aerial and satellite images, respectively. is the number of images we have used for each view in the training phase. is used to promote local semantic consistency and invariance between and , and defined as follows,
@@
InfoNCE Loss: In the PEE branch, although diversity and hierarchical embedding generation have been achieved, there exists a limitation of the consistency and discriminative capability. To address this issue, we introduce the InfoNCE loss to optimize the embedding space. By contrasting embeddings from different scenes (negative samples) and embeddings from the same scene (positive samples), InfoNCE loss ensures the consistency of embedding features within the same scene while maximizing the discriminability between embeddings of different scenes. The InfoNCE loss contains two queries from both detections, one of which uses a known satellite image to query aerial images, namely satellite drone. The other uses a known aerial image to query satellite images, which can be termed as dronesatellite. The InfoNCE Loss is defined as follows,
@@
CE Loss: In the GEE branch, the CE loss is introduced to optimize the global feature embeddings described in Eq.(5). The primary purpose of this loss function is to ensure that the model not only extracts rich semantic information from the global context, but also accurately maps these global feature embeddings to the correct class labels, thereby improving the classification performance. Through CE loss, the model learns the correspondence between global features and class labels, minimizing the discrepancy between the predicted and true distributions. CE loss measures the difference between the predicted class distribution of the global feature embeddings and the true class label. The loss function is defined as follows,
@@
Totle Loss: In the overall model, we combine three loss functions , , and . We optimize the overall performance of the model by minimizing these three functions. The total loss is formulated as follows,
@@ IV-A Experimental Datasets and Evaluation Metrics
To evaluate the performance of our proposed cross-view geo-localization framework, we conduct experiments on two large-scale datasets and one multi-weather scenario, including University-1652 [1], SUES-200 [35], and Multi-weather University-1652 [36]. These datasets provide complementary challenges for multi-view image matching and retrieval tasks.
University-1652 is a cross-view geo-localization dataset consisting of drone, satellite, and ground-level images from 1,652 locations across 72 universities worldwide. The training set includes 701 buildings from 33 universities, while the test set consists of 951 buildings from 39 universities, with no overlap between training and testing data. This dataset is first used to introduce drone view to cross-view geo-localization, extended the visual localization task as ground-drone-satellite cross-view matching.
SUES-200 introduces altitude variation in aerial images, featuring 200 unique locations. The training and test sets contain 120 and 80 locations, respectively. Each location includes a satellite image and aerial images taken at four altitudes: 150m, 200m, 250m, and 300m, encompassing diverse environments such as parks, lakes, and buildings. This dataset evaluates the model’s capability for cross-view retrieval with altitude variations.
Multi-weather University-1652 is an extension of University-1652 with ten simulated weather conditions, including fog, rain, and snow. It provides a benchmark to test the model’s robustness under various environmental conditions.
@@
These datasets present unique challenges such as altitude variation (SUES-200), large-scale distractor sets (University-1652), and weather-induced feature distortion (Multi-weather University-1652) to facilitate a comprehensive assessment of cross-view geo-localization performance.
@@
| Model | Venue | Paramars (M) | GFLOPs | R@1 | AP | R@1 | AP |
| LCM[11] | RS’2020 | 52.66 | - | 66.65 | 70.82 | 79.89 | 65.38 |
@@ IV-C Comparison with State-of-the-art Methods
We compare the proposed MEAN method with several state-of-the-art methods to demonstrate the superiority of our method. The experimental results on the University1652 dataset are reported in Table I, while the results on the Multi-weather University-165 dataset in Table II. Additionally, the results pertaining to the SUES-200 dataset are shown in Tables III and IV.
Results on University-1652: As shown in Table I, MAEN achieves 93.55% R@1 and 94.53% AP in the DroneSatellite setting, and 96.01% R@1 and 92.08% AP in the SatelliteDrone setting. These results demonstrate the strong performance and generalization capability of MAEN in cross-view geo-localization. MAEN offers an efficient and lightweight solution, reducing model size by 62.17% and computational complexity by 70. 99% compared to the current state-of-the-art model DAC [19], while maintaining comparable performance. Furthermore, with minimal parameter count and computational complexity, MAEN outperforms other advanced methods, including MCCG [7] for multi-feature representation, MFJR [17] for multi-branch joint optimization, and Sample4Geo [5], which employed the contrastive optimization techniques. These findings highlight MAEN’s ability to achieve high accuracy while maintaining computational efficiency, making it an effective model for cross-view geo-localization.
@@
| Model | Venue | Paramars(M) | GFLOPs | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
| LCM[11] | RS’2021 | 52.66 | - | 43.42 | 49.65 | 49.42 | 55.91 | 57.47 | 60.31 | 60.43 | 65.78 |
@@
| Model | Venue | Paramars(M) | GFLOPs | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
| LCM[11] | RS’2021 | 52.66 | - | 57.50 | 38.11 | 68.75 | 49.19 | 72.50 | 47.94 | 75.00 | 59.36 |
@@
| CAMP[30] | TGRS’2024 | 91.40 | 90.24 | 96.25 | 93.69 | 97.50 | 96.76 | 98.75 | 98.10 | 100.00 | 98.85 |
| DAC[19] | TCSVT’2024 | 103.84 | 90.24 | 97.50 | 94.06 | 98.75 | 96.66 | 98.75 | 98.09 | 98.75 | 97.87 |
@@
| Model | Venue | Paramars(M) | GFLOPs | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
@@
| Model | Venue | Paramars(M) | GFLOPs | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
@@
Results on SUES-200: As shown in Table III, in the DroneSatellite setting, MEAN achieves R@1 scores of 95.50%, 98.38%, 98.95%, and 99.52% and AP scores of 96.46%, 98.72%, 99.17%, and 99.63% at different altitude settings (150m, 200m, 250m, 300m). Although MEAN narrowly falls behind other models at 150m in terms of R@1 and AP, it surpasses state-of-the-art models at the other three altitude levels, achieving the best performance. This demonstrates MEAN’s strong adaptability in higher-altitude domains, effectively preserving high-level semantic consistency and robust across different views. Similarly, as shown in Table IV, in the SatelliteDrone setting, MEAN achieves R@1 scores of 97.50%, 100.00%, 100.00%, and 100.00%, and AP scores of 94.75%, 97.09%, 98.28%, and 99.21% across these altitude levels. MEAN consistently achieves the best performance at all heights, further confirming its stability and robustness in extracting and matching cross-view image features, regardless of altitude variations. In general, MEAN demonstrates substantial advantages in both performance and efficiency. Compared with the state-of-the-art models such as Sample4Geo, DAC, and CAMP [30], MEAN achieves a significant reduction in parameter count and computational complexity while maintaining outstanding accuracy. Among 18 evaluation metrics, MEAN achieves the best results in 16 metrics and ranks second in the remaining two, highlighting its robust ability to learn feature consistency and invariance when addressing cross-view perspective differences and scale variations. In addition to its ability to learn consistent and invariant features between cross-view images, MEAN exhibits high computational efficiency, making it an effective solution for cross-view geo-localization in large-scale scenarios.
@@ IV-D Comparison with State-of-the-Art Methods on Cross-Domain Generalization Performance
In cross-view geo-localization, cross-domain adaptability is a critical metric for evaluating the generalization capability of a model, especially when the training and testing datasets exhibit significant differences. To evaluate the transferability of our proposed model, we conducted experiments by training on the University-1652 dataset and testing on the SUES-200 dataset.
As shown in Table V and Table VI, MEAN demonstrates exceptional cross-domain adaptability in both DroneSatellite setting and SatelliteDrone setting across varying altitudes (150m, 200m, 250m, and 300m). In the DroneSatellite setting (Table V), MEAN achieves competitive results at all altitudes. At 150m, it attains an R@1 of 81.73% and an AP of 87.72%, surpassing other methods. As altitude increases, MEAN maintains robust performance, achieving R@1/AP scores of 89.05%/91.00% at 200m, 92.13%/93.60% at 250m and 94.63%/95.76% at 300m. These results highlight MEAN’s strong generalization capability across varying domains. In the SatelliteDrone setting (Table VI), MEAN similarly exhibits superior generalization performance. At 150m, it achieves the highest R@1 of 91.25% and AP of 81.50%, maintaining high competitiveness at higher altitudes. At 200m, 250m and 300m, MEAN outperforms other models with R@1/AP scores of 96.25%/89.55, 95.00%/92.36% and 96.25%/94.32%, respectively, further validating its robustness in cross-view geo-localization. Compared with the state-of-the-art models such as Sample4Geo, DAC, and CAMP, MEAN achieves superior performance while significantly reducing model complexity. With only 36.50M parameters and 26.18 GFLOPs, MEAN achieves the best performance in 13 out of 18 evaluation metrics and ranks second in four metrics. These findings underscore MEAN’s ability to learn domain-invariant features and adapt effectively to unseen data domains, making it a highly efficient and accurate model for cross-view geo-localization under cross-domain conditions.
@@ IV-E Ablation Studies
For loss functions, the joint application of and further improves the performance on cross-domain tasks. With all components and losses integrated, MEAN achieves the optimal results: R@1 and AP scores of 93.55% and 94.53% in the DroneSatellite setting, and 96.01% and 92.08% in the SatelliteDrone setting. Compared to the baseline, MEAN improves R@1 and AP by 9.55% and 7.92% in the DroneSatellite setting, and by 3.79% and 9.28% in the SatelliteDrone setting. These results validate the effectiveness of each component in advancing cross-view geo-localization.

### v4

@@ III-D Cross-Domain Enhanced Alignment branch
In CVGL tasks, images captured from different platforms often exhibit significant geometric differences, resulting in substantial variations in feature distributions and posing challenges for cross-domain matching. Feature alignment has been proven to be effective in addressing this issue [46, 47, 48]. Inspired by [22], we designed a cross-domain enhanced alignment (CEA) branch that uses multi-level fusion and adaptive calibration strategies to dynamically adjust feature consistency within a shared latent space in high-dimensional embeddings. As shown in Fig. 2, after the initial feature extraction by the backbone, a transformation is applied to flatten the spatial dimensions into a single dimension to obtain reshaped features , which serve as the input features for the CEA branch. Here, , .
@@
To enhance the consistency of cross-domain features, we introduce an adaptive calibration strategy in the CEA branch. Specifically, a convolutional layer is employed to project the two-dimensional input features into a higher-dimensional space. Subsequently, the batch normalization operation and the ReLU activation function are used to adjust the feature distribution. This procedure obtains a high-dimensional feature representation . Then, another convolutional layer is used to map the high-dimensional features back to the low-dimensional space, followed by the dropout operation and the normalization operation to improve the generalization capability. This produces a more compact feature representation , which improves computational efficiency and can be expressed as follows:
@@
After obtaining compact low-dimensional features , a 1×1 convolutional layer is applied to further adjust the feature representation. We apply T-Softmax to achieve the temperature-scaled local features with temperature factor , highlighting the crucial distinctions of the features . This is followed by S-Softmax for global re-balancing to ensure that the features can maintain consistency and harmony across different levels or scales. The output generated by the adaptive calibration strategy is denoted as , which captures both local distinctions and global consistency.
@@
To integrate the original and calibrated features, multi-level fusion is employed, which enables the aggregation of complementary representations and enhances cross-domain feature consistency. Specifically, both the input feature and the features are transformed into a shared dimension of by the convolutional layers and , respectively. Once the features have been transformed into the same dimensional space, they are concatenated along the channel dimension, namely the concatenated features represented as follows:
@@
where . Then a convolutional layer is applied to the concatenated features to restore the channel size back to , achieving a more compact representation. Finally, the batch normalization and the ReLU activation function are employed to normalize and activate the feature distributions. The output of the CEA branch is computed as follows:
@@ III-E Multi-Loss Optimization
For the feature embeddings and generated by the CEA branch described in Eq.(9), cosine similarity is used to estimate spatial directional consistency, while mean square errors measure absolute differences in the components of the feature. Using cosine similarity and mean square errors, CDA loss effectively realizes the semantic consistency of aligned features at a global scale and maintains consistency in local details. The loss function can be formulated as follows:
where and are weighting coefficients to balance the contributions of cosine similarity and mean square errors . is designed to promote global semantic consistency and invariance across cross-view features and represented as follows:
where and , are denoted index of the aerial and satellite images, respectively. is the number of images we have used for each view in the training phase. is used to promote local semantic consistency and invariance between and , and defined as follows:
@@
InfoNCE Loss: In the PEE branch, although diversity and hierarchical embedding generation have been achieved, there exists a limitation of the consistency and discriminative capability. To address this issue, we introduce the InfoNCE loss to optimize the embedding space. By contrasting embeddings from different scenes (negative samples) and embeddings from the same scene (positive samples), InfoNCE loss ensures the consistency of embedding features within the same scene while maximizing the discriminability between embeddings of different scenes. The InfoNCE loss contains two queries from both detections, one of which uses a known satellite image to query aerial images, namely satellite drone. The other uses a known aerial image to query satellite images, which can be termed as dronesatellite. The InfoNCE Loss is defined as follows:
@@
CE Loss: In the GEE branch, the CE loss is introduced to optimize the global feature embeddings described in Eq.(5). The primary purpose of this loss function is to ensure that the model not only extracts rich semantic information from the global context, but also accurately maps these global feature embeddings to the correct class labels, thereby improving the classification performance. Through CE loss, the model learns the correspondence between global features and class labels, minimizing the discrepancy between the predicted and true distributions. CE loss measures the difference between the predicted class distribution of the global feature embeddings and the true class label. The loss function is defined as follows:
@@
Total Loss: In the overall model, we combine three loss functions , , and . We optimize the overall performance of the model by minimizing these three functions. The total loss is formulated as follows:
@@ IV-A Experimental Datasets and Evaluation Metrics
To evaluate the performance of our proposed CVGL framework, we conduct experiments on two large-scale datasets and one multi-weather scenario, including University-1652 [1], SUES-200 [41], and Multi-weather University-1652 [40]. These datasets provide complementary challenges for multi-view image matching and retrieval tasks.
University-1652 is a CVGL dataset consisting of drone, satellite, and ground-level images from 1,652 locations across 72 universities worldwide. The training set includes 701 buildings from 33 universities, while the test set consists of 951 buildings from 39 universities, with no overlap between training and testing data. This dataset is first used to introduce drone view to CVGL, extended the visual localization task as ground-drone-satellite cross-view matching.
SUES-200 introduces altitude variation in aerial images, featuring 200 unique locations. The training and test sets contain 120 and 80 locations, respectively. Each location includes a satellite image and aerial images captured at four altitudes, encompassing diverse environments such as parks, lakes, and buildings. This dataset evaluates the model’s capability for cross-view retrieval with altitude variations.
Multi-weather University-1652 is an extension of University-1652 with ten simulated weather conditions providing a benchmark for evaluating model robustness under diverse environmental scenarios.
@@
These datasets present unique challenges such as altitude variation (SUES-200), large-scale distractor sets (University-1652), and weather-induced feature distortion (Multi-weather University-1652) to facilitate a comprehensive assessment of CVGL performance.
@@
| Model | Venue | Parameters (M) | GFLOPs | R@1 | AP | R@1 | AP |
@@ IV-C Comparison with State-of-the-art Methods
We compare MEAN with state-of-the-art methods to demonstrate its effectiveness. The experimental results on the University-1652 dataset are reported in Table I, while the results on the Multi-weather University-165 dataset in Table II. Additionally, the results pertaining to the SUES-200 dataset are shown in Tables III and IV.
Results on University-1652: As shown in Table I, MAEN achieves 93.55% R@1 and 94.53% AP in the DroneSatellite setting, and 96.01% R@1 and 92.08% AP in the SatelliteDrone setting. These results demonstrate the strong performance and generalization capability of MAEN in CVGL. MAEN offers an efficient and lightweight solution, reducing model size by 62.17% and computational complexity by 70. 99% compared with the current state-of-the-art model DAC [22], while maintaining comparable performance. Furthermore, with minimal parameter count and computational complexity, MAEN outperforms other advanced methods, including MCCG [10] for multi-feature representation, MFJR [20] for multi-branch joint optimization, and Sample4Geo [5], which employed the contrastive optimization techniques. These findings highlight MAEN’s ability to achieve high accuracy while maintaining computational efficiency, making it an effective model for CVGL.
@@
| Model | Venue | Parameters(M) | GFLOPs | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
@@
| Model | Venue | Parameters(M) | GFLOPs | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
@@
| Sample4Geo[5] | ICCV’2023 | 87.57 | 90.24 | 97.50 | 93.63 | 98.75 | 96.70 | 98.75 | 98.28 | 98.75 | 98.05 |
| DAC[22] | TCSVT’2024 | 96.50 | 90.24 | 97.50 | 94.06 | 98.75 | 96.66 | 98.75 | 98.09 | 98.75 | 97.87 |
@@
| Model | Venue | Parameters(M) | GFLOPs | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
@@
| Model | Venue | Parameters(M) | GFLOPs | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
@@
Results on SUES-200: As shown in Table III, in the DroneSatellite setting, MEAN achieves R@1 scores of 95.50%, 98.38%, 98.95%, and 99.52% and AP scores of 96.46%, 98.72%, 99.17%, and 99.63% at different altitude settings (150m, 200m, 250m, 300m). Although MEAN narrowly falls behind other models at 150m in terms of R@1 and AP, it surpasses state-of-the-art models at the other three altitude levels, achieving the best performance. This demonstrates MEAN’s strong adaptability in higher-altitude domains, effectively preserving high-level semantic consistency and robust across different views. Similarly, as shown in Table IV, in the SatelliteDrone setting, MEAN achieves R@1 scores of 97.50%, 100.00%, 100.00%, and 100.00%, and AP scores of 94.75%, 97.09%, 98.28%, and 99.21% across these altitude levels. MEAN consistently achieves the best performance at all heights, further confirming its stability and robustness in extracting and matching cross-view image features, regardless of altitude variations. In general, MEAN demonstrates substantial advantages in both performance and efficiency. Compared with the state-of-the-art models such as Sample4Geo, DAC, MEAN achieves a significant reduction in parameter count and computational complexity while maintaining outstanding accuracy. Among 18 evaluation metrics, MEAN achieves the best results in 16 metrics and ranks second in the remaining two, highlighting its robust ability to learn feature consistency and invariance when addressing cross-view perspective differences and scale variations. In addition to its ability to learn consistent and invariant features between cross-view images, MEAN exhibits high computational efficiency, making it an effective solution for CVGL in large-scale scenarios.
@@ IV-D Comparison with State-of-the-Art Methods on Cross-Domain Generalization Performance
In CVGL, cross-domain adaptability is a critical metric for evaluating the generalization capability of a model, especially when the training and testing datasets exhibit significant differences. To evaluate the transferability of our proposed model, we conducted experiments by training on the University-1652 dataset and testing on the SUES-200 dataset.
As shown in Table V and Table VI, MEAN demonstrates exceptional cross-domain adaptability in both DroneSatellite setting and SatelliteDrone setting across varying altitudes (150m, 200m, 250m, and 300m). In the DroneSatellite setting (Table V), MEAN achieves competitive results at all altitudes. At 150m, it attains an R@1 of 81.73% and an AP of 87.72%, surpassing other methods. As altitude increases, MEAN maintains robust performance, achieving R@1/AP scores of 89.05%/91.00% at 200m, 92.13%/93.60% at 250m and 94.63%/95.76% at 300m. These results highlight MEAN’s strong generalization capability across varying domains. In the SatelliteDrone setting (Table VI), MEAN similarly exhibits superior generalization performance. At 150m, it achieves the highest R@1 of 91.25% and AP of 81.50%, maintaining high competitiveness at higher altitudes. At 200m, 250m and 300m, MEAN outperforms other models with R@1/AP scores of 96.25%/89.55, 95.00%/92.36% and 96.25%/94.32%, respectively, further validating its robustness in CVGL. Compared with the state-of-the-art models such as Sample4Geo, DAC, MEAN achieves superior performance while significantly reducing model complexity. With only 36.50M parameters and 26.18 GFLOPs, MEAN achieves the best performance in 13 out of 18 evaluation metrics and ranks second in four metrics. These findings underscore MEAN’s ability to learn domain-invariant features and adapt effectively to unseen data domains, making it a highly efficient and accurate model for CVGL under cross-domain conditions.
@@ IV-E Ablation Studies
For loss functions, the joint application of and further improves the performance on cross-domain tasks. With all components and losses integrated, MEAN achieves the optimal results: R@1 and AP scores of 93.55% and 94.53% in the DroneSatellite setting, and 96.01% and 92.08% in the SatelliteDrone setting. Compared with the baseline, MEAN improves R@1 and AP by 9.55% and 7.92% in the DroneSatellite setting, and by 3.79% and 9.28% in the SatelliteDrone setting. These results validate the effectiveness of each component.
| Setting (DEG) | University-1652 | |||||
| DroneSatellite | SatelliteDrone | |||||
| R@1 | AP | R@1 | AP | |||
| 93.09 | 94.19 | 95.86 | 92.72 | |||
| 92.24 | 93.46 | 95.29 | 91.96 | |||
| 92.61 | 93.75 | 95.72 | 92.38 | |||
| 93.32 | 94.33 | 95.44 | 92.51 | |||
| 93.14 | 94.16 | 96.14 | 92.40 | |||
| 92.41 | 93.61 | 96.00 | 91.98 | |||
| 93.16 | 94.23 | 95.86 | 92.76 | |||
| 92.87 | 93.94 | 95.29 | 92.28 | |||
| 92.64 | 93.77 | 95.44 | 92.14 | |||
| 93.55 | 94.54 | 96.01 | 92.08 | |||
| Setting (DEC) | University-1652 | |||||
| DroneSatellite | SatelliteDrone | |||||
| R@1 | AP | R@1 | AP | |||
| 93.24 | 94.30 | 95.43 | 92.55 | |||
| 93.32 | 94.38 | 95.72 | 92.76 | |||
| 93.29 | 94.36 | 95.15 | 92.62 | |||
| 92.92 | 94.06 | 95.29 | 92.56 | |||
| 93.19 | 94.31 | 96.15 | 92.80 | |||
| 93.18 | 94.29 | 95.58 | 92.54 | |||
| 93.11 | 94.20 | 95.58 | 92.27 | |||
| 92.81 | 93.93 | 95.29 | 92.71 | |||
| 93.04 | 94.12 | 95.72 | 92.76 | |||
| 93.55 | 94.54 | 96.01 | 92.08 | |||
To further investigate the effectiveness of using different dilation rates in the DEG and DEC modules, the dilation rates have been adjusted to various configurations. As shown in Table LABEL:Ablation_Studies1, when the dilation rates of the DEC module are fixed at 1,2,3, the experimental performance demonstrates that uniform configurations with same dilation rates in DEG module generally lead to suboptimal performance compared with the configurations with various dilation rates. From this table, the progressive pattern with dilation rates 1,2,3 in DEG module achieves the better results. Similarly, it can be observed that the progressive pattern with dilation rates 1,2,3 in DEC module also achieves the better results when the dilation rates of the DEG module are fixed at 1,2,3 in Table LABEL:Ablation_Studies2. Furthermore, we conduct the experiment using the same dilation rates in both DEG and DEC modules as shown in Table LABEL:Ablation_Studies3. The results demonstrate that when both modules share the varied dilation rate combinations, the model outperforms configurations with a single fixed dilation rate. Moreover, setting both modules to 1,2,3 achieves the best overall results.
| Setting (DEG+DEC) | University-1652 | |||||
| DroneSatellite | SatelliteDrone | |||||
| R@1 | AP | R@1 | AP | |||
| 93.44 | 94.47 | 95.15 | 92.58 | |||
| 93.20 | 94.31 | 95.95 | 92.90 | |||
| 92.60 | 93.80 | 95.86 | 91.96 | |||
| 93.48 | 94.49 | 95.72 | 92.73 | |||
| 93.12 | 94.20 | 96.14 | 92.65 | |||
| 93.26 | 94.30 | 95.72 | 92.52 | |||
| 93.14 | 94.22 | 95.01 | 92.21 | |||
| 92.84 | 93.99 | 95.29 | 92.28 | |||
| 93.29 | 94.38 | 95.29 | 92.32 | |||
| 93.55 | 94.54 | 96.01 | 92.08 | |||